QUESTION 9

9.1 State De Moivre’s Theorem.
(2)

9.2 Express cos 5θ and sin 4θ as polynomials in terms of sin θ and cos θ.
(8)

9.3 Let w be a negative real number, z a 6th root of w.

(a) Show that z (k) = ρ 6

6th roots of w.

cos

( π+2kπ )

+ i sin

( π+2kπ )

, k = 0, 1, 2, 3, 4, 5 is a formula for the



Show all your working.
(8)

(b) Hence determine the 6th roots of −729.
(2)

(c) Given z = cos θ + i sin θ and u + iv = (1 + z)(1 + z2). Prove that v = u tan( 3θ ) and

u2 + v2 = 16 cos2( θ ) cos2(θ)
(10)

[30]
1
Expert's answer
2020-05-04T20:09:56-0400

9.1 The Moivre’s Theorem

"z=r(\cos\theta+i\sin \theta)\\
z^n=r^n(\cos n\theta+i\sin n\theta), n\in N"


9.2 Let "r=1"

"z=\cos\theta+i\sin\theta\\
z^5=\cos5\theta+i\sin5\theta=(\cos\theta+i\sin\theta)^5=\\
=\cos^5\theta+5i\cos^4\theta\sin\theta+10i^2\cos^3\theta\sin^2\theta+\\
+10i^3\cos^2\theta\sin^3\theta+5i^4\cos\theta\sin^4\theta+i^5\sin^5\theta=\\
=\cos^5\theta-10\cos^3\theta\sin^2\theta+5\cos\theta\sin^4\theta+\\
+i(5\cos^4\theta\sin\theta-10\cos^2\theta\sin^3\theta+\sin^5\theta)\\
\cos5\theta=\cos^5\theta-10\cos^3\theta\sin^2\theta+5\cos\theta\sin^4\theta\\"


"z^4=\cos4\theta+i\sin4\theta=(\cos\theta+i\sin\theta)^4=\\
=\cos^4\theta+4i\cos^3\theta\sin\theta+6i^2\cos^2\theta\sin^2\theta+\\
+6i^3\cos\theta\sin^3\theta+i^4\sin^4\theta=\\
=\cos^4\theta-6\cos^2\theta\sin^2\theta+\sin^4\theta+\\
+i(4\cos^3\theta\sin\theta-6\cos\theta\sin^3\theta)\\
\sin4\theta=4\cos^3\theta\sin\theta-6\cos\theta\sin^3\theta"


"i^2=-1, i^3=-i, i^4=1,i^5=i"


9.3 a)

"z=\sqrt[6]{w}, w<0\\
w=\rho(\cos\theta+i\sin\theta)\\
z=\sqrt[6]\rho(\cos\frac{\theta+2\pi k}{6}+i\sin\frac{\theta+2\pi k}{6}), k=0,1,2,3,4,5"


b)

"\sqrt[6]{-729}=\sqrt[6]{-1\cdot3^6}=3\sqrt[6]{-1}\\
-1=\cos\pi+i\sin\pi\\
\sqrt[6]{-1}=\cos\frac{\pi+2\pi k}{6}+\sin\frac{\pi+2\pi k}{6}, k=0,1,2,3,4,5\\
\sqrt[6]{-729}=3(\cos\frac{\pi+2\pi k}{6}+\sin\frac{\pi+2\pi k}{6}),\\
 k=0,1,2,3,4,5\\"


"\sqrt[6]{-729}=3(\cos\frac{\pi}{6}+\sin\frac{\pi}{6})=3(\frac{\sqrt3}{2}+i\frac{1}{2}),\\
\sqrt[6]{-729}=3(\cos\frac{3\pi }{6}+\sin\frac{3\pi }{6})=3i,\\
\sqrt[6]{-729}=3(\cos\frac{5\pi }{6}+\sin\frac{5\pi }{6})=3(-\frac{\sqrt3}{2}+i\frac{1}{2}),\\
\sqrt[6]{-729}=3(\cos\frac{7\pi}{6}+\sin\frac{7\pi}{6})=3(-\frac{\sqrt3}{2}-i\frac{1}{2}),\\
\sqrt[6]{-729}=3(\cos\frac{9\pi}{6}+\sin\frac{9\pi}{6})=-3i,\\
\sqrt[6]{-729}=3(\cos\frac{11\pi }{6}+\sin\frac{11\pi }{6})=3(\frac{\sqrt3}{2}-i\frac{1}{2})\\"


c)

"z=\cos\theta+i\sin\theta\\
u+iv=(1+z)(1+z^2)=\\
=((1+\cos\theta)+i\sin\theta)\cdot\\(1+\cos^2\theta+2i\cos\theta\sin\theta-\sin^2\theta)=\\
=((1+\cos\theta)+i\sin\theta)\cdot\\((1+\cos^2\theta-\sin^2\theta)+2i\cos\theta\sin\theta)=\\
=((1+\cos\theta)+i\sin\theta)\cdot\\(2\cos^2\theta+2i\cos\theta\sin\theta)=\\
=2\cos^2\theta(1+\cos\theta)-2\cos\theta\sin^2\theta+\\
+i(2\cos\theta\sin\theta(1+\cos\theta)+2\sin\theta\cos^2\theta)\\
u=2\cos^2\theta(1+\cos\theta)-2\cos\theta\sin^2\theta=\\
=2\cos\theta(\cos\theta+\cos2\theta)\\
v=2\cos\theta\sin\theta(1+\cos\theta)+2\sin\theta\cos^2\theta=\\
=2\sin\theta\cos\theta(1+2\cos\theta)=2\cos\theta(\sin\theta+\sin2\theta)\\
u\tan\frac{3\theta}{2}=2\cos\theta(\cos\theta+\cos2\theta)\frac{\sin\frac{3\theta}{2}}{\cos\frac{3\theta}{2}}=\\
=\frac{2\cos\theta}{\cos\frac{3\theta}{2}}\frac{\sin\frac{\theta}{2}+\sin\frac{5\theta}{2}-\sin\frac{\theta}{2}+\sin\frac{7\theta}{2}}{2}=\\
=\frac{2\cos\theta}{\cos\frac{3\theta}{2}}\frac{2\sin3\theta\cos\frac{\theta}{2}}{2}=\\
=2\cos\theta(2\sin\frac{3\theta}{2}\cos\frac{\theta}{2})=\\
=2\cos\theta(\sin\theta+\sin2\theta)=v"


"u^2+v^2=u=4\cos^2\theta(\cos\theta+\cos2\theta)^2+\\
+4\cos^2\theta(\sin\theta+\sin2\theta)^2=\\
=4\cos^2\theta(\cos^2\theta+2\cos\theta\cos2\theta+\cos^22\theta+\\
+\sin^2\theta+2\sin\theta\sin2\theta+\sin^22\theta)=\\
=4\cos^2\theta(2+2\cos\theta)=8\cos^2\theta(1+\cos\theta)=\\
=16\cos^2\theta\cos^2\frac{\theta}{2}"